# Test-time Computing: from System-1 Thinking to System-2 Thinking

### v1

@@ 4.2.2 Self-correction
Most of the aforementioned self-correction methods demonstrate significant performance improvements on advanced closed-source large models or open-source LLMs with over 70B parameters. However, for medium-scale open-source models with weaker capabilities, we need to further fine-tune them to unlock their self-correction capabilities. Supervised fine-tuning optimizes the model using high-quality multi-turn correction data, either manually annotated (Saunders et al., 2022) or sampled from stronger LLMs (Paul et al., 2024; Qu et al., 2024; Gao et al., 2024b; Zhang et al., 2024g; Xi et al., 2024). GLoRe (Havrilla et al., 2024) considers that LLMs need global or local refinement for different types of errors. To address this, they construct training sets for global and local refinement, train verifiers to identify global and local errors, and develop LLMs for refinement based on different global or local feedback signals. Although SFT is effective, training data from offline-generated self-correction trajectories can only simulate limited correction patterns. This leads to the distribution mismatch with the actual self-correction behavior during model inference. Self-correct (Welleck et al., 2023) adopts online imitation learning, re-sampling new self-correction trajectories for training after each training epoch. SCoRe (Kumar et al., 2024) proposes using the multi-turn RL method to improve self-critique and self-correction capability.
@@ 4.2.3 Tree Searching
Repeated sampling and self-correction scale test-time computation in parallel and sequentially, respectively. Human thinking is a tree search that combines brainstorming in parallel with backtracking to find other paths to solutions when it encounters a dead end. Search algorithms and value functions are two critical components in tree searching.
@@
In LLM reasoning, current search algorithms include uninformed search and heuristic search. Uninformed search does not rely on specific heuristic information but explores the search space according to a fixed rule. For example, tree-of-thought (ToT) (Yao et al., 2023) adopts the BFS or DFS to search, while Xie et al. (2023) use beam search. Uninformed search is usually less efficient for problems with large search spaces, so heuristic search represented by MCTS is widely used in reasoning tasks (Hao et al., 2023; Zhang et al., 2024b; Bi et al., 2024). MCTS gradually optimizes search results through four steps: selection, expansion, simulation, and backpropagation, approaching the optimal solution. Long (2023) trains an LLM controller using reinforcement learning to guide the LLM reasoner’s search path.
@@
The value function evaluates the value of each action and guides the tree to expand towards branches with higher values in MCTS. RAP (Hao et al., 2023) designs a series of heuristic value functions, including the likelihood of the action, the confidence of the state, self-evaluation results, and task-specific reward, and combines them according to task requirements. Reliable and generalized value functions facilitate the application of MCTS to more complex problems with deeper search spaces. AlphaMath (Chen et al., 2024a) and TS-LLM (Feng et al., 2024b) replace the hand-crafted value function with a learned LLM value function, automatically generating reasoning process and step-level evaluation signals in MCTS. Traditional MCTS methods expand only one trajectory, while rStar (Qi et al., 2024) argues that the current value function struggles to guide the selection of the optimal path accurately. Therefore, rStar retains multiple candidate paths and performs reasoning with another LLM, ultimately selecting the path where both LLMs’ reasoning results are consistent. Gao et al. (2024c) propose SC-MCTS inspired by contrast decoding, which utilizes multiple external reward models as value functions.
@@
Tree search can guide LLMs to generate long reasoning processes, and these data help train LLMs with stronger reasoning abilities. ReST-MCTS* (Zhang et al., 2024a) uses process rewards as a value function to guide MCTS, collecting high-quality reasoning trajectories and the value of each step to improve the policy model and reward model. Due to the step-by-step exploration of tree search, it can obtain finer-grained step-level feedback signals. MCTS-DPO (Xie et al., 2024) collects step-level preference data through MCTS and uses DPO for preference learning. Recently, many o1-like technical reports (Qin et al., 2024b; Zhao et al., 2024b; Zhang et al., 2024f) have also confirmed the necessity of using tree search to construct high-quality long reasoning chain data for training.
@@ 5.1 Generalizable System-2 Model
Currently, most o1-like models exhibit strong deep reasoning abilities only in specific domains such as math and code and struggle to adapt to cross-domain or general tasks. The key to addressing this issue lies in enhancing the generalization ability of verifiers or critics. Currently, some works utilize model ensemble (Lin et al., 2024b) or regularization constraints (Yang et al., 2024; Jia, 2024) to make verifiers more generalizable. Nevertheless, there is still significant room for improvement in the generalization of the verifier. Additionally, weak-to-strong generalization (Burns et al., 2023) is a topic worth further exploration. People are no longer satisfied with solving mathematical problems with standard answers; they hope System-2 models can assist in scientific discovery and the proofs of mathematical conjectures. In such cases, even human experts struggle to provide accurate feedback, while weak-to-strong generalization offers a promising direction to address this issue.
@@ 5.2 Multimodal Reasoning
In System-1 thinking, TTA has been successfully applied to multimodal LLMs, improving performance in tasks such as zero-shot image classification, image-text retrieval, and image captioning (Zhao et al., 2024a). However, test-time computing methods in System-2 thinking remain limited to text modalities. Visual, speech, and other modalities are crucial for model understanding and interaction with the world. To achieve cognitive intelligence, System-2 models must be able to fully integrate multimodal information for reasoning. The exploration of multimodal CoT (Zhang et al., 2024h; Wu et al., 2024b; Mondal et al., 2024; Lee et al., 2024; Gao et al., 2024a) opens up the possibility of building multimodal System-2 models. Xu et al. (2024) are the first to apply test-time computing to visual reasoning tasks. They divide the visual reasoning process into four stages: task summary, caption, reasoning, and answer conclusion. They propose a stage-level beam search method, which repeatedly samples at each stage and selects the best result to the next stage. We believe test-time computing still holds significant potential for development in multimodal reasoning. For example, incorporating more modalities like speech and video into reasoning tasks, applying successful methods such as reflection mechanisms and tree search to multimodal reasoning, or aligning the multimodal reasoning process with human cognitive processes.
@@ 5.3 Efficiency and Performance Trade-off
The successful application of test-time computing shows that sacrificing reasoning efficiency can lead to better reasoning performance. However, researchers continue to seek a balance between performance and efficiency, aiming to achieve optimal performance under a fixed reasoning latency budget. This requires adaptively allocating computational resources for each sample. Damani et al. (2024) train a lightweight module to predict the difficulty of a question, and allocate computational resources according to its difficulty. Zhang et al. (2024c) further extend the allocation targets to more hyperparameters. There are still many open questions worth exploring, such as how to integrate inference acceleration strategies, e.g. KV cache compression, token pruning, and speculative decoding with test-time computing and how to predict problem difficulty more accurately.
@@ 5.4 Scaling Law
Unlike training-time computation scaling, test-time computing still lacks a universal scaling law. Some works have attempted to derive scaling laws for specific test-time computing strategies (Wu et al., 2024c; Chen et al., 2024d). Brown et al. (2024) demonstrate that the performance has an approximately log-linear relationship with repeated sampling times. Snell et al. (2024) investigate the scaling laws of repeated sampling and self-correction, and propose the computing-optimal scaling strategy. There are two major challenges to achieving a universal scaling law: first, current test-time computation strategies are various, each with different mechanisms to steer the model; thus, it lacks a universal framework for describing them; second, the performance of test-time computation is affected by a variety of factors, including the difficulty of samples, the accuracy of feedback signals, and decoding hyperparameters, and we need empirical studies to filter out the critical factors.
@@ 5.5 Strategy Combination
Different test-time computing strategies are suited to various tasks and scenarios, so combining multiple strategies is one way to achieve better System-2 thinking. For example, Marco-o1 (Zhao et al., 2024b) combines the MCTS and self-correction, using MCTS to plan reasoning processes, and self-correction to improve the accuracy of each step. Moreover, test-time adaptation strategies in System-1 models can also be combined with test-time reasoning strategies. Akyürek et al. (2024) combine test-time training with repeated sampling. They further optimize the language modeling loss on test samples, then generate multiple candidate answers through data augmentation, and finally determine the answer by majority voting. They demonstrate the potential of test-time training in reasoning tasks, surpassing the human average on the ARC challenge. Therefore, we think that for LLM reasoning, it is crucial to focus not only on emerging test-time strategies but also on test-time adaptation methods. By effectively combining these strategies, we can develop System-2 models that achieve or surpass o1-level performance.

### v2

@@ 4.2.2 Self-correction
Most of the aforementioned self-correction methods demonstrate significant performance improvements on advanced closed-source large models or open-source LLMs with over 70B parameters. However, for medium-scale open-source models with weaker capabilities, we need to further fine-tune them to unlock their self-correction capabilities. Supervised fine-tuning optimizes the model using high-quality multi-turn correction data, either manually annotated (Saunders et al., 2022), self-rationalize (Zelikman et al., 2022; Yuan et al., 2025b), multi-agent debate (Subramaniam et al., 2025) or sampled from stronger LLMs (An et al., 2023; Paul et al., 2024; Qu et al., 2024; Gao et al., 2024c; Zhang et al., 2024k; Xi et al., 2024). GLoRe (Havrilla et al., 2024) considers that LLMs need global or local refinement for different types of errors. To address this, they construct training sets for global and local refinement, train verifiers to identify global and local errors, and develop LLMs for refinement based on different global or local feedback signals. Xi et al. (2024) design a scalable framework for synthesizing self-correction training data, enabling reasoning models to generate controlled errors and receive feedback from critics to self-correct. Although SFT is effective, training data from offline-generated self-correction trajectories can only simulate limited correction patterns. This leads to the distribution mismatch with the actual self-correction behavior during model inference. Self-correct (Welleck et al., 2023) adopts online imitation learning, re-sampling new self-correction trajectories for training after each training epoch. To further expand the exploration space of LLMs, many studies adopt flexible RL algorithms to surpass the performance limits of SFT. SCoRe (Kumar et al., 2024) proposes using the multi-turn RL method to improve self-critique and self-correction capability. T1 (Hou et al., 2025) employs self-correction training data for SFT cold-start, followed by RL training using the RLOO algorithm (Ahmadian et al., 2024). During the RL phase, high-temperature sampling and entropy rewards encourage the LLM to explore more diverse reasoning paths. Deepseek-R1 (Guo et al., 2025a) uses rule-based rewards and the GRPO algorithm (Shao et al., 2024) for RL training. It also demonstrates RL’s immense potential, even without SFT cold-start, its exploration capabilities suffice to endow LLMs with strong reasoning abilities.
@@ 4.2.3 Tree Searching
Repeated sampling and self-correction scale test-time compute in parallel and sequentially, respectively. Human thinking is a tree search that combines brainstorming in parallel with backtracking to find other paths to solutions when it encounters a dead end. Search algorithms and value functions are two critical components in tree searching.
@@
In LLM reasoning, current search algorithms include uninformed search and heuristic search. Uninformed search explores the search space according to a fixed rule. For example, tree-of-thought (ToT) (Yao et al., 2023) adopts the BFS or DFS to search, while Xie et al. (2023) use beam search. Uninformed search is usually less efficient for problems with large search spaces, so heuristic search strategies represented by A (Meng et al., 2024; Wang et al., 2024a) and MCTS (Hao et al., 2023; Bi et al., 2024; Park et al., 2024) are widely used in reasoning tasks. MCTS, which eliminates the need for explicit heuristics, leverages stochastic simulations and adaptive tree expansion under uncertain environments, making it well-suited for large state spaces. It optimizes search results gradually through four steps: selection, expansion, simulation, and backpropagation, approaching the optimal solution. In contrast, A uses a heuristic function-guided deterministic search to guarantee optimal paths, but its performance depends on the design of the heuristic function. As a result, MCTS has been successfully applied to tasks such as RAG (Hu et al., 2024b; Jiang et al., 2024; Li and Ng, 2024; Feng et al., 2025), QA (Luo et al., 2025a; Gan et al., 2025), hallucinations mitigation (Cheng et al., 2025), text-to-SQL (Yuan et al., 2025a), etc. Additionally, Long (2023) trains an LLM controller using reinforcement learning to guide the LLM reasoner’s search path, and Chari et al. (2025) utilizes ant colony evolutionary algorithm to guide tree search.
@@
The value function evaluates the value of each state and guides the tree to expand towards branches with higher values in heuristic tree search. Xu (2023) train an energy function by noise-contrastive estimation as the value function. RAP (Hao et al., 2023) designs a series of heuristic value functions, including the likelihood of the action, the confidence of the state, self-evaluation results, and task-specific reward, and combines them according to task requirements. Reliable and generalized value functions facilitate the application of MCTS to more complex problems with deeper search spaces. AlphaMath (Chen et al., 2024a) and TS-LLM (Feng et al., 2024b) replace the hand-crafted value function with a learned LLM value function, automatically generating reasoning process and step-level evaluation signals in MCTS. VerifierQ (Qi et al., 2024a) integrates implicit Q-learning and contrastive Q-learning to train the value function, effectively mitigating the overestimation issue at the step level. Traditional MCTS methods expand only one trajectory, while rStar (Qi et al., 2024b) argues that the current value function struggles to guide the selection of the optimal path accurately. Therefore, rStar retains multiple candidate paths and performs reasoning with another LLM, ultimately selecting the path where both LLMs’ reasoning results are consistent. Gao et al. (2024d) propose SC-MCTS which combines multiple reward models, including contrastive reward, likelihood and self-evaluation as value functions. MCTSr (Zhang et al., 2024b) and SR-MCTS (Zhang et al., 2024c) take complete responses as nodes, expanding the search space through self-critique and correction. SR-MCTS utilizes pairwise preference rewards and global quantile score as the value function, offering more robust value function estimation compared to step-based MCTS.
@@
Tree search can guide LLMs to generate long reasoning processes, and these data help train LLMs with stronger reasoning abilities (Zhai et al., 2024; Xu et al., 2024a; Guan et al., 2025). ReST-MCTS* (Zhang et al., 2024a) uses process rewards as a value function to guide MCTS, collecting high-quality reasoning trajectories and the value of each step to improve the policy model and reward model. Due to the step-by-step exploration of tree search, it can obtain finer-grained step-level feedback signals. MCTS-DPO (Xie et al., 2024) collects step-level preference data through MCTS and uses DPO for preference learning. AlphaLLM-CPL (Wang et al., 2024k) ranks trajectories based on preference reward gaps and policy prediction gaps, employing curriculum learning to efficiently utilize MCTS-collected trajectories. Recently, many o1-like technical reports (Qin et al., 2024b; Zhao et al., 2024b; Zhang et al., 2024j) have also confirmed the necessity of using tree search to construct high-quality long reasoning chain data for training.
Summary 2:
Repeated sampling is easy to implement and improves answer diversity, making it suitable for open-ended or easily verifiable tasks, though computationally inefficient. Self-correction relies on precise, fine-grained feedback and works well for easily verifiable tasks, but may not perform well with poor feedback or weak reasoning capability. Tree search optimizes complex planning tasks globally but involves complex implementation.
@@ 5.1 Generalizable System-2 Model
Currently, most o1-like models exhibit strong reasoning abilities only in specific domains such as math and code and struggle to adapt to cross-domain or general tasks. The key to addressing this issue lies in enhancing the generalization ability of verifiers or critics (LeVine et al., 2023; Kim et al., 2024d; Chen et al., 2024c). Currently, some works utilize multi-objective training (Wang et al., 2024c), model ensemble (Lin et al., 2024b) or regularization constraints (Yang et al., 2024a; Jia, 2024) to make verifiers more generalizable. Nevertheless, there is still significant room for improvement in the generalization of the verifier. Additionally, weak-to-strong generalization (Burns et al., 2023) is a topic worth further exploration. People are no longer satisfied with solving mathematical problems with standard answers; they hope System-2 models can assist in scientific discovery and the proofs of mathematical conjectures. In such cases, even human experts struggle to provide accurate feedback, while weak-to-strong generalization offers a promising direction to address this issue (Tang et al., 2025). We think that more generalized System-2 models may not rely on a single feedback source but instead obtain multi-source feedback through interactions between LLM-based agents and tools, experts, or other agents (Nathani et al., 2023; Lan et al., 2024).
@@ 5.2 Multimodal Reasoning
In System-1 thinking, TTA has been successfully applied to multimodal LLMs, improving performance in tasks such as zero-shot image classification, image-text retrieval, and image captioning (Zhao et al., 2024a). However, test-time compute methods in System-2 thinking remain limited to text modalities. Visual, speech, and other modalities are crucial for model understanding and interaction with the world. To achieve cognitive intelligence, System-2 models must be able to fully integrate multimodal information for reasoning. The exploration of multimodal CoT (Zhang et al., 2024l; Wu et al., 2024b; Mondal et al., 2024; Lee et al., 2024; Gao et al., 2024b) and multimodal critics or verifiers (Xiong et al., 2024) open up the possibility of building multimodal System-2 models. Xu et al. (2024b) are the first to apply test-time compute to visual reasoning tasks. They divide the visual reasoning process into four stages: task summary, caption, reasoning, and answer conclusion. They propose a stage-level beam search method, which repeatedly samples at each stage and selects the best result for the next stage. Nowadays, Qwen team has released the open-weight multimodal reasoning model QVQ (Qwen, 2024), OpenAI and Kimi (Team et al., 2025) have released their multimodal reasoning products. We believe test-time compute still holds significant potential for development in multimodal reasoning. For example, incorporating more modalities like speech and video into reasoning tasks, applying successful methods such as reflection mechanisms and tree search (Yao et al., 2024; Dong et al., 2024a) to multimodal reasoning, or aligning the multimodal reasoning process with human cognitive processes. Besides understanding and reasoning tasks, Xie et al. (2025) and Guo et al. (2025b) show test-time compute can improve image generation performance, with great potential for multimodal generation in the future.
@@ 5.3 Efficiency and Performance Trade-off
The successful application of test-time compute shows that sacrificing reasoning efficiency can lead to better reasoning performance. However, researchers continue to seek a balance between performance and efficiency, aiming to achieve optimal performance under a fixed reasoning latency budget. This requires adaptively allocating computational resources for each sample. Damani et al. (2024) train a lightweight module to predict the difficulty of a question, and allocate computational resources according to its difficulty. Zhang et al. (2024f) further extend the allocation targets to more hyperparameters. Chen et al. (2025b) and Wang et al. (2025) systematically evaluate the over-thinking and under-thinking phenomena in o1-like models, where the former leads models to overcomplicate simple problems, and the latter causes frequent switching of reasoning paths on difficult problems, thereby reducing reasoning efficiency. O1-Pruner (Luo et al., 2025b) propose the length-penalty PPO loss to shorten reasoning processes while maintaining accuracy. There are still many open questions worth exploring, such as how to integrate inference acceleration strategies, e.g. model compression (Li et al., 2024f; Huang et al., 2024c; Li et al., 2025b), token pruning (Fu et al., 2024; Zhang et al., 2024d), and speculative decoding (Leviathan et al., 2023; Xia et al., 2024) with test-time compute, and how to allocate optimal reasoning budget according to problem difficulty.
@@ 5.4 Scaling Law
Unlike training-time computation scaling, test-time compute still lacks a universal scaling law. Some works have attempted to derive scaling laws for specific test-time compute strategies (Wu et al., 2024c; Levi, 2024). Brown et al. (2024) demonstrate that the performance has an approximately log-linear relationship with repeated sampling times. Chen et al. (2024e) models repeated sampling as a knockout tournament and league-style algorithm, proving theoretically that the failure probability of repeated sampling follows a power-law scaling. Snell et al. (2024) investigate the scaling laws of repeated sampling and self-correction, and propose the computing-optimal scaling strategy. There are two major challenges to achieving a universal scaling law: first, current test-time compute strategies are various, each with different mechanisms to steer the model; thus, it lacks a universal framework for describing them; second, the performance of test-time compute is affected by a variety of factors, including the difficulty of samples, the accuracy of feedback signals, and decoding hyperparameters, and we need empirical studies to filter out the critical factors.
@@ 5.5 Strategy Combination
Different test-time compute strategies are suited to various tasks and scenarios, so combining multiple strategies is one way to achieve better System-2 thinking. For example, Marco-o1 (Zhao et al., 2024b) combines the MCTS and self-correction, using MCTS to plan reasoning processes, and self-correction to improve the accuracy of each step. TPO (Li et al., 2025a) combines BoN sampling and self-correction. Moreover, test-time adaptation strategies in System-1 models can also be combined with test-time reasoning strategies. Akyürek et al. (2024) combine test-time training with repeated sampling. They further optimize the language modeling loss on test samples, then generate multiple candidate answers through data augmentation, and finally determine the answer by majority voting. They demonstrate the potential of test-time training in reasoning tasks, surpassing the human average on the ARC challenge. Therefore, we think that for LLM reasoning, it is crucial to focus not only on emerging test-time strategies but also on test-time adaptation methods. By effectively combining these strategies, we can develop System-2 models that achieve or surpass o1-level performance.
6 Benchmarks and Open-source Frameworks
6.1 Benchmarks
Test-time Adaptation
In System-1 models, distribution shifts include adversarial robustness, cross-domain and cross-lingual scenarios. In the field of CV, ImageNet-C (Hendrycks and Dietterich, 2019), ImageNet-R (Hendrycks et al., 2021a), ImageNet-Sketch (Wang et al., 2019) are common datasets for TTA. Yu et al. (2023) propose a benchmark to conduct a unified evaluation of TTA methods across different TTA settings and backbones on 5 image classification datasets. For NLP tasks, TTA is primarily applied in QA and machine translation tasks, with commonly used datasets such as MLQA (Lewis et al., 2020), XQuAD (Artetxe et al., 2020), MRQA (Fisch et al., 2019), CCMatrix (Schwenk et al., 2021) and Ted Talks (Qi et al., 2018).
Feedback Modeling
RewardBench (Lambert et al., 2024) collects 20.2k prompt-choice-rejection triplets covering tasks such as dialogue, reasoning, and safety. It evaluates the accuracy of reward models in distinguishing between chosen and rejected responses. RM-Bench (Liu et al., 2024d) further evaluates the impact of response style on reward models. RMB (Zhou et al., 2024) extends the evaluation to the more practical BoN setting, where reward models are required to select the best response from multiple candidates. CriticBench (Lin et al., 2024c) is specifically designed to evaluate a critic model’s generation, critique, and correction capabilities. For PRM, Song et al. (2025a) propose PRMBench, which evaluates PRMs whether they can identify the earliest incorrect reasoning step in math tasks. ProcessBench (Zheng et al., 2024) provides a more fine-grained evaluation, including redundancy, soundness, and sensitivity. In addition, there are benchmarks for evaluating multimodal feedback modeling, such as VL-RewardBench (Li et al., 2024d) and MJ-Bench (Chen et al., 2024f).
Test-time Reasoning
Reasoning capability is the core of System-2 models, including mathematics, code, commonsense, planning, etc (Zeng et al., 2024). Math reasoning is one of the most compelling reasoning tasks. With the advancements in LLM and test-time compute, the accuracy on some previously challenging benchmarks, like GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b), have surpassed the 90% mark. Thus, more difficult college admissions exam (Zhang et al., 2023b; Arora et al., 2023; Azerbayev et al., 2024) and competition-level (Gao et al., 2024a) math benchmarks have been proposed. Some competition-level benchmarks are not limited to textual modalities in algebra, logic reasoning, and word problems. For instance, OlympiadBench (He et al., 2024), OlympicArena (Huang et al., 2024b) and AIME (Zamil and Rabby, 2024) provide images for geometry problems, incorporating visual information to aid in problem-solving, while AlphaGeometry (Trinh et al., 2024) employs symbolic rules for geometric proofs. The most challenging benchmark currently is FrontierMath (Glazer et al., 2024), with problems crafted by mathematicians and covering major branches of modern mathematics. Even the most advanced o3 has not achieved 30% accuracy.
Code ability is a key aspect of LLM reasoning, with high practical value, covering code completion (Ding et al., 2023; Zhang et al., 2023a; Gong et al., 2024a), code reasoning (Gu et al., 2024), and code generation (Chen et al., 2021; Austin et al., 2021) tasks. Among these, code generation gains more attention. HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) provide natural language descriptions of programming problems, requiring LLMs to generate corresponding Python code and use unit tests for evaluation. MultiPL-E (Cassano et al., 2022) extend them to 18 program languages. EvalPlus (Liu et al., 2024a) automatically augments test cases to assess the robustness of the generated code. Recently, some studies collect benchmarks from open-source projects, which are closed to realistic applications and more challenging due to complex function calls, such as DS-1000 (Lai et al., 2023), CoderEval (Yu et al., 2024a), EvoCodeBench (Li et al., 2024a) and BigCodeBench (Zhuo et al., 2025).
Commonsense reasoning requires LLMs to possess both commonsense knowledge and reasoning abilities. Early benchmarks (Zellers et al., 2019; Talmor et al., 2019; Sakaguchi et al., 2021; Bisk et al., 2020) focus on evaluating LLMs’ commonsense ability. StrategyQA (Geva et al., 2021) collects more complex and subtle multi-hop reasoning questions. MMLU (Hendrycks et al., 2021b) and MMLU-Pro (Wang et al., 2024n) cover commonsense reasoning questions across various domains, including STEM, the humanities, the social sciences, etc. Planning aims to enable LLMs to take optimal actions based on the current state and environment to complete tasks. Current planning benchmarks primarily focus on synthetic tasks, such as Blocksworld (Valmeekam et al., 2023b), Crosswords, and Game-of-24 (Yao et al., 2023).
6.2 Projects
OpenR
(Wang et al., 2024d) [^2] https://github.com/openreasoner/openr is an open-source test-time reasoning framework that integrates various test-time compute strategies, PRM training, and improvement training. It currently supports beam search, BoN, MCTS, and rStar, and implements popular online reinforcement learning algorithms like APPO, GRPO, and TPPO.
RLHFlow
(Dong et al., 2024b) offers a comprehensive framework for reward modeling [^3] https://github.com/RLHFlow/RLHF-Reward-Modeling and online RLHF training [^4] https://github.com/RLHFlow/Online-RLHF. Its standout feature is the integration of various reward model training methods, including the vanilla preference reward model, multi-objective reward models, PRM, etc.
OpenRLHF
(Hu et al., 2024a) [^5] https://github.com/OpenRLHF/OpenRLHF also integrates reward modeling and RLHF training but focuses more on the efficient implementation of reinforcement learning algorithms and training tricks. Its strength lies in the integration of distributed training and efficient fine-tuning, enabling users to easily train large language models with more than 70B parameters.